# A review on combinatorial approach to aggregation

### v1

@@ Abstract
Recently, a combinatorial approach to discrete, finite, and irreversibly aggregating systems has been progressively developed. In this work, we review its achievements up to the present moment, focusing on the practical aspects and discussing its limitations. First, we present the assumptions and combinatorial foundations of the approach, which are based on direct counting of the system states, in contrast to the previous approaches of Smoluchowski and Marcus–Lushnikov. A method to obtain combinatorial expressions for the average number of clusters of a given size and the corresponding standard deviation is described by solving the simplest example of a constant kernel. Then, we extend consideration to a number of kernels (e.g., additive, product, linear–chain, condensation), which were recently solved by explicitly finding the number of internal states of the cluster of a given size. Next, we show that theoretical predictions for any given kernel may be obtained with no need to find an explicit solution but using a recursive expression. We exploit this opportunity to present the use of combinatorial expressions to solve kernels related to the real processes of aerosol growth and planetesimal formation. At this point, a comparison to numerical results appears. Other potential application fields are indicated, including dust agglomeration and polymer growth. Finally, issues related to the varying precision of the theoretical predictions are summarized. In the last section, we propose open problems.
@@
Figure 1: Blood coagulation, or clotting, is the process by which blood changes from a liquid to a gel, forming a blood clot. This image shows a scanning electron micrograph of a number of red blood cells found enmeshed in a fibrinous matrix on the luminal surface of an indwelling vascular catheter. Reproduced from a public domain wiki_commons_2005 .
@@ I Introduction
$\left(i\right)+\left(j\right){{\stackrel{{\scriptstyle K\left(i,j\right)}}{{\longrightarrow}}}}\left(i+j\right),$ (1)
where $\left(i\right)$ stands for a cluster of property $i$ and $K\left(i,j\right)$ is the coagulation kernel representing the rate of the process. Property $i$ is usually regarded as mass or size but may also be regarded as other additive quantities such as length or electric charge or (this is often the case in real processes) as an effective quantity considering all related properties of the particle. Since the system is closed and the clusters grow in time, all particles eventually join to form a single giant cluster.
@@
Figures 1 and 2 present examples of processes from different fields where particles undergo clustering. In Fig. 1, physiological blood clotting is depicted wiki_commons_2005 . Fig. 2 illustrates two snapshots from a system of drying aerosol Frohlich_2023 .
@@
The first to study aggregation was Smoluchowski Smoluchowski_1917 . Kinetic differential equations of his approach require several assumptions characteristic of the mean–field theory, such as the infinite volume of the system and continuous cluster concentrations. They were intensively studied for decades and are widely used nowadays. However, a limited number of aggregation processes have explicit solutions to the Smoluchowski equation. It has been used to derive analytical solutions for a few particular kernels, e.g., constant, $K(i,j)=\mathrm{const}$, additive, $K(i,j)\propto i+j$, and multiplicative (product), $K(i,j)\propto ij$, and for selected initial conditions. Other kernels may be studied with approximations, numerically, or using the so–called scaling solutions. Several extensions beyond aggregation itself are known, e.g., aggregation–fragmentation or aggregation with mass injection Blum_2006 ; Leyvraz_2003 ; Krapivsky_2010 ; Banasiak_2019 .
The next successful attempt at analytical description was the stochastic approach by Marcus Marcus_1968 , later developed by Lushnikov Lushnikov_2004 ; Lushnikov_2005 ; Lushnikov_2011 . Within this approach, exact solutions of the master equations for the basic kernels were obtained for the simplest case of monodisperse initial conditions, as well as for the mixture of monomers and dimers Fronczak_2019 . The basic kernels were also revisited in a thermodynamic–focused description using the so–called linear ensembles 2018_Matsoukas .
In a recent decade, a novel (combinatorial) approach has been proposed, where probability distributions were obtained through direct counting of the system states grassberger1 ; grassberger2 ; Fronczak_2018 ; Lepek_2019 . Recently, this approach has proved its usefulness in solving more complex forms of kernels with arbitrary parameters under monodisperse initial conditions, i.e., for the linear–chain kernel Lepek_2021 , $K\propto(1/i+1/j)^{\alpha}$, condensation kernel, $K\propto(A+j)(A+j)$, and the sum of constant and additive kernels Lepek_2021_ROMP , $K\propto A+i+j$. For these kernels, explicit expressions for the average number of clusters of a given size and corresponding standard deviation were obtained. The accuracy of those theoretical solutions tested against numerical simulation and experimental data varied from approximate to exact depending on the parameter and coagulation time. By the term “exact”, we mean solutions to the coagulation process in contrast to the exact solutions of the Smoluchowski coagulation equation, which, in general, is not the same.
This review will expose that the combinatorial approach goes further than the abovementioned kernel forms and may be easily used to obtain theoretical predictions for any given kernel. To do so, we produced auxiliary numerical results of planetesimal formation and aerosol growth to compare them to the theoretical predictions. Moreover, we present the importance of the ability to predict standard deviation from the mean together with theoretical solutions (using the example of planetesimals). We propose and briefly outline several practical fields where this method may be helpful: polymer growth, protein aggregation, coagulation of soot, milk, and magnetorheological fluids, cosmic matter aggregation in planetary rings, and dust agglomeration.
Throughout this work, we will relate to tests of the theoretical predictions against numerical simulations. A comment on that must be given here. In general, an aggregation process may be studied with several traditional methods: direct numerical simulation of the system, solving the Smoluchowski equation analytically, solving the Smoluchowski equation numerically, or by other (indirect) methods such as scaling solutions. Still, even the numerical solution may be challenging, as the system of equations is usually infinite. Monte Carlo methods have been recently used as a possible way to deal with this problem Osinsky_2022 . Recently, semi-analytical methods have also been used to study aggregation (still aimed at solving Smoluchowski equation) Kaushik_2023 .
@@
Authors sometimes refer to numerical simulations, meaning, for instance, numerical solutions to the Smoluchowski equation. In our work, when we refer to numerical simulation, we always mean direct numerical simulation, where the current state of the system is held in the computer memory and is updated step by step strictly due to the kernel rule. Such a simulation may be regarded as ground truth for any analytical approach as it simulates the aggregation process itself, giving insight into any property of the system without any “proxy” tool. Details on how to implement such a ground-truth simulation for any given kernel may be found, for instance, in Lepek_2021 . For the analytically–derived expressions (both explicit or recursive), we use the terms theoretical solution or theoretical prediction.
@@
The review is built as follows. In Section II, we thoroughly described the initial assumptions and construction of combinatorial expressions. Also, a primary tool to obtain explicit theoretical solutions (a generating function method) has been presented. Section III presents a review of known explicit theoretical solutions to several aggregation kernels. Section IV shows that no explicit solution is needed to obtain theoretical predictions if a recursive expression may be used. Two examples of use for two different aggregation processes are presented to give a better picture of usability. Further fields of potential use are listed. Readers not keenly interested in the mathematical foundations of the approach may find this section illuminating. Sections V and VI summarize the approach and discuss its limitations. Open problems are proposed in the last section. Mathematical tools (partial Bell polynomials and Lagrange inversion) required to understand derivations and transformations are outlined in corresponding sections in the Appendix.
@@ II.1 Initial assumptions
In the combinatorial methodology to analyze the aggregating system, we assume: (i) relatively low particle concentrations, since we do not consider the possibility of having encounters among three or more particles, or several encounters at a given time, (ii) monodisperse initial conditions (in the initial state, all of the clusters are monomers of size one), (iii) individual coagulation acts are instantaneous and take place in continuous time, and (iv) discrete-time labels which count subsequent coagulation acts, i.e., subsequent states of the system.
The system consists of $N$ monomers (the smallest units, monomer particles), some of which later join into clusters. $N=\text{const}$ is equivalent to the preservation of mass in the system. A single coagulation act occurs in one time step, therefore, the total number of clusters, $k$, at time step $t$ is
@@
where $n_{s}\geq 0$ stands for the number of clusters of mass $s$ (where $s$ is the number of monomeric units included in the cluster) and $n_{1}$ corresponds to monomers, $n_{2}$ to dimers, $n_{3}$ to trimers, and so on. This results in the following system state in the initial moment of the process:
@@
$\sum^{N}_{s=1}{n_{s}=k}\;\;\;\;\textrm{and}\;\;\;\;\sum^{N}_{s=1}{sn_{s}=N}.$ (6)
The above basic assumptions are analogous to the Marcus–Lushnikov approach; however, in the combinatorial approach, calculations include direct counting of available states instead of using the Master equation.
II.2 Remark on discrete vs. physical time
@@
Time labels mark subsequent coagulation acts (time steps). The first coagulation act occurs at time step $t=1$, the second at time step $t=2$, and so on. These discrete time labels are a counter to subsequent states of the system.
@@
In a real coagulating system, particles merge (through binary collisions) with frequency determined by the kernel and by the evolution stage of the system. However, the exact time between these acts remains unknown (it is stochastic). It is modeled in, e.g., Marcus–Lushnikov framework and may be modeled also in the combinatorial approach as its further extension. However, it must be emphasized that the discrete timescale of binary aggregation acts used throughout this work is not the same as the physical time measured, e.g., in seconds. Binary aggregation time labels may be regarded as a (nonlinearly) rescaled version of the physical time and do not provide instantaneous information on the reaction rate expressed in collisions per second.
@@
However, it is crucial to emphasize that the lack of direct relevance to the physical time is not a severe limitation. Suppose we have information on the initial number of clusters and a “snapshot” of the cluster size distribution at a particular point in time. In that case, the translation to the language of discrete time steps is straightforward because a particular number of clusters in the system clearly defines the time step in the combinatorial formalism. Such a procedure of comparing combinatorial predictions to the experimental data has been shown in Ref. Lepek_2021 .
Figure 3: Possible ways to group a set of $N=5$ elements into two clusters of sizes 2 and 3. All particles are distinguishable. There are ten ways to do so (cf. Eq. (8)). Illustration by the authors.

### v2

# A mini-review on combinatorial solutions to the Marcus–Lushnikov irreversible aggregation

@@ Abstract
Over the last decade, a combinatorial approach to discrete, finite, and irreversibly aggregating systems has been progressively developed. In this work, we review its achievements up to the present moment, focusing on the practical aspects and discussing its limitations. First, we present the assumptions and combinatorial foundations of the approach, which are based on direct counting of the system states, in contrast to the previous approaches of Smoluchowski and Marcus–Lushnikov. A method to obtain combinatorial expressions for the average number of clusters of a given size and, importantly, the corresponding standard deviation is described by solving the simplest example of a constant kernel. An expression for a complete probability distribution for a number of clusters of a given size is also presented. Then, we extend consideration to a number of kernels (e.g., additive, product, linear-chain, condensation), which were recently solved by explicitly finding the number of ways to create a cluster of a given size. We show that, for a general case, the present framework yields approximate solutions. In this way, theoretical predictions for any given kernel may be obtained with no need to find an explicit solution but using a recursive expression. We exploit this opportunity to present the use of combinatorial expressions to solve kernels related to the real processes of aerosol growth and planetesimal formation. At this point, a comparison to numerical results appears. Finally, issues related to the validity and varying precision of the theoretical predictions are summarized. In the last section, we propose open problems. Appendix contains partial Bell polynomials, generating function method, Lagrange inversion, potential fields of further application, and considerations on the relation of the presented combinatorial solutions to the scaling solutions of the Smoluchowski equation.
@@
Figure 1: Blood coagulation, or clotting, is the process by which blood changes from a liquid to a gel, forming a blood clot. This image shows a scanning electron micrograph of a number of red blood cells found enmeshed in a fibrinous matrix on the luminal surface of an indwelling vascular catheter. Reproduced from the public domain wiki_commons_2005 .
@@ I Introduction
$\left(s\right)+\left(g\right){{\stackrel{{\scriptstyle K\left(s,g\right)}}{{\longrightarrow}}}}\left(s+g\right),$ (1)
where $\left(s\right)$ and $(g)$ stand for clusters of property $s$ and $g$, respectively, and $K\left(s,g\right)$ is the coagulation kernel representing the rate of the process. Property $s$ is usually regarded as mass or size but may also be regarded as other additive quantities, such as length or electric charge, or (this is often the case in real processes) as an effective quantity considering all related properties of the particle. Since the system is closed and the clusters grow in time, all particles eventually join to form a single giant cluster.
@@
Figs. 1 and 2 present examples of processes from different fields where particles undergo clustering. In Fig. 1, physiological blood clotting is depicted wiki_commons_2005 . Fig. 2 illustrates two snapshots from a system of drying aerosol Frohlich_2023 .
The first to study aggregation was Smoluchowski Smoluchowski_1917 . Kinetic differential equations of his approach require several assumptions characteristic of the mean-field theory, such as the infinite number of particles, continuous cluster concentrations, and perfect mixing of the system (spatial invariance). They were intensively studied for decades and are widely used nowadays. However, a limited number of aggregation processes have explicit solutions to the Smoluchowski equation. It has been used to derive analytical solutions for the three classic kernels, i.e., the constant, $K(s,g)=\mathrm{const}$, additive, $K(s,g)\propto s+g$, and multiplicative (product), $K(s,g)\propto sg$, and for selected initial conditions. The above solutions were found for both continuous and discrete versions (here, continuous and discrete relate to the allowed sizes of particles). A variety of other kernels (see, e.g., Refs. daCosta_1998 ; Aldous_1999 ) may be studied within this approach with approximations, numerically, or using the so-called scaling solutions. Several extensions are known, e.g., aggregation–fragmentation, aggregation with mass injection Blum_2006 ; Leyvraz_2003 ; Krapivsky_2010 ; Banasiak_2019 , or the so-called truncated Smoluchowski equations used in numerical solutions and finite-mass systems daCosta_1998 . As such, they find their applications (see, e.g., Rouwhorst_2020_Nature ). However, due to the nature of the Smoluchowski formalism, one can ask several fundamental questions. How large does the system have to be for a correct comparison to the results from the Smoluchowski equation? What (obvious) discrepancies are expected to arise between the Smoluchowski theory and the ground truth aggregation process of the finite coagulating system after the initial stage, when the Smoluchowski assumptions are not met? The above questions, together with the limited number of explicit solutions known, motivate the search for new approaches.
The next successful attempt at analytical description of the irreversible aggregation was the stochastic approach by Marcus Marcus_1968 , later studied and developed by Lushnikov Lushnikov_1978 ; Lushnikov_2004 ; Lushnikov_2005 ; Lushnikov_2011 . It may be regarded as a microscopic model of the deterministic Smoluchowski model, both describing spatially averaged systems (spatially invariant) and being quantitatively similar for sufficiently large systems. However, the Marcus–Lushnikov approach gives a more detailed description of the process than Smoluchowski. It allows us to ask questions—such as the probability distribution of particles of given size at a given time or the variability between realizations—that are beyond the scope of the deterministic Smoluchowski rate equations vanDongen_1987_fluct ; vanDongen_1987_fluct_II ; Leyvraz_2022 . Another key difference concerns large clusters (large with respect to the system size). While the Smoluchowski equation allows for clusters of unbounded size and represents averages over all possible orders of aggregation events, the Marcus–Lushnikov model represents one possible sequence of coagulation events and restricts clusters to size less than a finite number (finite system size, NN). This makes it possible to study the distributions of clusters for any given size, including those similar to NN in both non-gelling and gelling systems Leyvraz_2022 . In general, this stochastic Marcus–Lushnikov model provides expected time-dependent size-distribution profiles (either by analytical solutions or averaging over many numerical simulations), which can be compared to solutions of the Smoluchowski model. It also gives information on the distribution of possible resulting states of the system, which has no equivalent in the Smoluchowski approach (cf. Sec. II.9).
Within the Marcus-Lushnikov approach, exact solutions of the master equation for the basic kernels (constant, additive, product) were obtained for the simplest case of monodisperse initial conditions, as well as for the mixture of monomers and dimers Fronczak_2019 . Mathematical complexity of this approach prevented further extension to other kernels.
Later, the basic kernels were also revisited in a thermodynamic-focused description using the so-called linear ensembles 2018_Matsoukas .
In a recent decade, a novel (combinatorial) approach has been proposed to further overcome the shortcomings of the Smoluchowski approach by providing detailed information on the cluster size distribution in finite systems, even for the late stage of their evolution, when the continuous cluster concentration conditions are no longer valid. This discrete model is based on the Marcus–Lushnikov assumptions, but, in this case, probability distributions were obtained through direct counting of the system states for the three basic kernels grassberger1 ; grassberger2 ; Fronczak_2018 ; Lepek_2019 . Recently, this approach has proved its usefulness in solving more complex forms of kernels with arbitrary parameters under monodisperse initial conditions, i.e., for the linear-chain kernel Lepek_2021 , $K\propto(1/s+1/g)^{\alpha}$, condensation kernel, $K\propto(A+s)(A+g)$, and the sum of constant and additive kernels Lepek_2021_ROMP , $K\propto A+s+g$. For these kernels, explicit expressions for the average number of clusters of a given size and corresponding standard deviation were obtained.
The assumptions behind the work on the combinatorial approach were only achieved to a certain extent: in general, the solutions found were only approximate. The accuracy of those combinatorial theoretical solutions tested against numerical simulation and experimental data varied from approximate to exact depending on the parameters and coagulation time. By the term “exact”, we mean solutions to the finite-system coagulation process in contrast to the exact solutions of the Smoluchowski coagulation equation, which, in general, is not the same. As we will see, the relation of the asymptotic behavior of the combinatorial solutions to the known results from the Smoluchowski approach (scaling theory and similarity solutions) presents a complex view Leyvraz_2022 , with the asymptotic forms of the combinatorial solutions agreeing closely with the Smoluchowski solutions for the constant and the additive kernels, while differing, especially for small clusters, for the general case.
Despite the limited number of works in the area of the combinatorial solutions, they address a variety of different aggregation processes and provide a tool to make an attempt to model an arbitrary process, which was not explicitly described before. Therefore, we believe that a concise report of the topic in the form of this (mini-)review article, providing advantages, limitations, and connections to the usual Smoluchowski equations, may be valued in several fields, including finite-size experimental setups.
At this point, we will also signal another key difference between the Smoluchowski framework and the combinatorial (discrete Marcus–Lushnikov) framework, which relates to defining the time; continuous in the Smoluchowski equation and discrete in the latter case. Although a large-scale and long-time transformation between these two timescales may be used, comparing combinatorial solutions to the Smoluchowski solutions needs utmost caution. This issue is described in detail in Section II.2. Another consequence of the discrete time of the combinatorial approach is its characteristic insensitivity to multiplying kernel forms by a constant factor. This is why we usually use the proportionality sign, $\propto$”, when defining kernels in this work. Implications are discussed in Sec. II.2.
This mini-review will expose that the combinatorial approach may be easily used to obtain theoretical predictions of the average number of clusters of a given size for any given kernel, although yielding, in general, only approximate results. To do so, we produced auxiliary numerical results of planetesimal formation and aerosol growth to compare them to the theoretical predictions. We show the importance of the ability to predict standard deviation from the mean, together with theoretical solutions (using the example of planetesimals). We propose and briefly outline several practical fields where this method may be helpful: polymer growth, protein aggregation, coagulation of soot, milk, and magnetorheological fluids, cosmic matter aggregation in planetary rings, and dust agglomeration.
Moreover, besides presenting an insight into the first moment (average number of clusters) and the second moment (standard deviation) of the time-dependent probability distribution of finding a number of clusters of a given size, we show a method to find a complete cluster size distribution.
Throughout this work, we will relate to tests of the theoretical predictions against numerical simulations. A comment on that must be given here. In general, an aggregation process may be studied with several traditional methods: direct numerical simulation of the system, solving the Smoluchowski equation analytically, solving the Smoluchowski equation numerically, or using other (indirect) methods such as scaling solutions. Still, even the numerical solution may be challenging, as the system of equations is usually infinite. Monte Carlo methods have been recently used as a possible way to deal with this problem Osinsky_2022 . Recently, semi-analytical methods have also been used to study aggregation (still aimed at solving the Smoluchowski equation) Kaushik_2023 .
@@
Authors sometimes refer to numerical simulations, meaning, for instance, numerical solutions to the Smoluchowski equation. In our work, when we refer to numerical simulation, we always mean direct numerical simulation of a system with perfect mixing, where the current state of the system is held in the computer memory and is updated step by step with binary aggregation acts strictly due to the kernel rule. It is done by assigning probabilities to each possible aggregation act and by randomly selecting one. Such a simulation may be regarded as reference for the analytical Marcus–Lushnikov approach, as it simulates the aggregation process itself, giving insight into any property of the system without any “proxy” tool. Details on how to implement such a numerical simulation for any given kernel may be found, for instance, in Refs. Kang_1986 ; Lepek_2021 . Throughout this work, we use the implementation from Ref. Lepek_2021 . For the analytically-derived expressions (both explicit or recursive), we use the terms theoretical solution or theoretical prediction. For the cases for which the precision of the analytical solution is already determined, we clearly specify whether it is an exact solution or an approximation. For the cases that are still challenging in such determination, we give an appropriate reference to the section where the details are considered.
At this moment, to clearly outline the key contributions and methodological framework discussed in this mini-review, we may summarize the main points as follows. An exact combinatorial solution for the constant kernel is presented, expressed in terms of Bell polynomials. Building on this, an ansatz is introduced to extend the solution to arbitrary kernels. This extension results in an uncontrolled approximation in all cases other than the constant (and possibly the additive) kernel. Nonetheless, the approximation remains qualitatively correct for large particles. In addition, for a large variety of kernels, this approximate solution can still be evaluated in closed form.
@@
The mini-review is built as follows. In Section II, we thoroughly described the initial assumptions and construction of combinatorial expressions. Also, a primary tool to obtain explicit theoretical solutions (a generating function method) has been presented. Derivations of expressions are shown for an average number of clusters of a given size, corresponding standard deviation, and probability distribution of finding a given number of clusters of a given size. Section III presents a review of known explicit theoretical solutions to several aggregation kernels. Section IV shows that no explicit solution is needed to obtain theoretical predictions (approximations for a general case) if a recursive expression may be used. Two examples of use for two different aggregation processes are presented to give a better picture of usability. Readers not keenly interested in the mathematical foundations of the approach may find this section illuminating. Sections V and VI summarize the approach and discuss its limitations. In particular, in Section V, we provide a brief discussion on how combinatorial solutions relate to the scaling solutions of the Smoluchowski equation. Open problems are proposed in the last section. Mathematical tools (partial Bell polynomials and Lagrange inversion) required to understand derivations and transformations are outlined in the corresponding sections in the Appendix. Also, further fields of potential use are listed. In Appendix C, we provide an introduction to the scaling theory of aggregating systems and present related large-size and large-time limit of a general combinatorial solution as compared to the known results from the scaling theory for several kernels. The last part of the Appendix contains usual Smoluchowski solutions to the three basic kernels.
@@ II.1 Initial assumptions
In the combinatorial methodology to analyze the aggregating system, we assume: (i) relatively low particle concentrations, since we do not consider the possibility of having encounters among three or more particles, or several encounters at a given time, (ii) monodisperse initial conditions (in the initial state, all of the clusters are monomers of size 1), (iii) individual coagulation acts are instantaneous and take place in continuous time, and (iv) discrete-time labels which count subsequent coagulation acts, i.e., subsequent states of the system.
The system consists of NN monomers (the smallest units, monomer particles), some of which later join into clusters. $N=\text{const}$ is equivalent to the preservation of mass in the system. A single coagulation act occurs in one time step; therefore, the total number of clusters, $k$, at time step $t$ is
@@
where $n_{s}\geq 0$ stands for the number of clusters of size $s$ (where $s$ is the number of monomeric units included in the cluster) and $n_{1}$ corresponds to monomers, $n_{2}$ to dimers, $n_{3}$ to trimers, $n_{4}$ to tetramers, and so on.
$\sum^{N}_{s=1}{n_{s}=k}\;\;\;\;\textrm{and}\;\;\;\;\sum^{N}_{s=1}{sn_{s}=N}.$ (4)
The above basic assumptions, Eqs. (3) and (4), are analogous to the original Marcus–Lushnikov approach.
They result in the following system state in the initial moment of the process:
@@
For the next state $\Omega(2)$, there are two possible coagulation steps: two monomers can join to create a second dimer with relative probability $K(1,1)$, or a monomer and a dimer can join to create a trimer with relative probability $K(1,2)$. In this way, probabilities of the next states, $\Omega(2)$, $\Omega(3)$, and so on, are determined by the kernel $K$ applied.
Finding probabilities $P(\Omega)$ for any given time and system is, in general, a non-trivial task. In the original Marcus–Lushnikov approach, they were found by solving the master equation (cf., e.g., Ref. Fronczak_2019 ). Instead, in the combinatorial approach presented here, finding these probabilities will include direct counting of available system states.
II.2 Discrete vs. continuous time. Relation to Smoluchowski formalism
@@
Time labels mark subsequent coagulation acts (time steps). The first coagulation act occurs at time step $t=1$, the second at time step $t=2$, and so on. These discrete time labels are a counter to subsequent states of the system. The last aggregation act occurs in t=N1t=N-1 (thus, for large systems, $t\approx N$).
This is a key difference of the timescale in this model as compared to the continuous-time models of Smoluchowski and the continuous-time (original) Marcus–Lushnikov model Marcus_1968 ; Lushnikov_1978 . A continuous Marcus–Lushnikov time variable equivalent to the discrete combinatorial timescale $t$ may be simply found as $\theta=t/N$. Therefore, at a fixed time point in the continuous-time Marcus–Lushnikov model, the number of particles is a random variable, whereas for the present discrete (combinatorial) model, the number of particles (clusters) is given at each time step Leyvraz_2022 .
With this formulation, a transformation may be found to connect the time in the combinatorial approach, $t$, to the time in the Smoluchowski approach, $\tilde{t}$. As t=Nkt=N-k is the discrete time in the combinatorial approach, and if $c_{s}(\tilde{t})=n_{s}/N$ corresponds to concentrations of clusters of size $s$, then
$\sum_{s=1}^{\infty}c_{s}(\tilde{t})=1-\frac{t}{N},$ (7)
provides a correct relation between $\tilde{t}$ of the Smoluchowski equations and $t$ of the discrete Marcus–Lushnikov model (cf. Eqs. (165) and (166) in Appendix C) Leyvraz_2022 .
Using the above relation, it can be shown that the distribution of the cluster numbers $z=n_{s}/N$ as a function of $t$ tends towards $)\delta(z-c_{s}(\tilde{t}$)). As a result, the only difference between the two frameworks is that the Smoluchowski theory provides a less detailed description than the Marcus–Lushnikov theory, as the former does not describe the full distribution of cluster numbers, and always involves a limit $N\to\infty$. In this way, for large values of NN, the distribution of $n_{s}$ is sharply peaked around the value given by the Smoluchowski theory. However, for small NN, for which the distribution does not reduce to a delta function, Marcus–Lushnikov theory provides a fuller view. What also follows is that the variance of the distribution is beyond the reach of the Smoluchowski theory, though it vanishes for the limit of $N\to\infty$ Leyvraz_2022 .
For a formally correct comparison between the above two approaches, one shall find infinite-size approximations of the solutions to the discrete combinatorial model as in Ref. Leyvraz_2022 (note that for a particular point in the aggregation time, $\theta$, assuming $N\rightarrow\infty$ implies $t\rightarrow\infty$ as well). For the introduction to the similarity solutions and their relation to the large-scale limits of the combinatorial solutions, please see Appendix C.
At this point, an experimental researcher will raise the following concern. Obviously, both combinatorial solutions and Smoluchowski solutions undergo comparison to the experimental data, and comparing predictions from these two frameworks for finite systems is a logical eventuality. Therefore, for practical purposes, in this work, we will plot the solutions to the Smoluchowski equation compared to the combinatorial solutions and to the result from numerical simulations for the constant, additive, and product kernels. For these three basic kernels, the Smoluchowski equation has been solved exactly, yielding explicit (closed form) solutions. We find empirically the time $\tilde{t}$ for the Smoluchowski solutions that correspond to the discrete time $t$ using Eq. (7). For complete results of this analysis, please see Fig. 9.
@@
In a real-time coagulating system, particles merge (through binary collisions) with frequency determined by the kernel and by the evolution stage of the system (i.e., concentrations). However, the exact time between these acts remains unknown (it is stochastic in nature). It was modeled previously in Marcus–Lushnikov framework Eibeck_2000 ; Fronczak_2019 and may be modeled also in the combinatorial approach as its further extension. However, as we know from the above considerations, the discrete timescale of binary aggregation acts (as well as $\theta$) used in the combinatorial formalism is not the same as the physical time measured, e.g., in seconds. Binary aggregation time labels may be regarded as a rescaled version of the physical time and do not provide instantaneous information on the reaction rate expressed in collisions per second. For the purposes of transformation, Eq. (7) may be used.
@@
However, at the cost of a somewhat complex relationship with real time, we obtain a different benefit. Suppose we have information on the initial number of clusters and a “snapshot” of the cluster size distribution at a particular point in time (thus, NN and $k$). In that case, the translation to the language of discrete time steps is straightforward because a particular number of clusters in the system clearly defines the time step in the combinatorial formalism (t=Nkt=N-k). Such an empirical procedure of comparing combinatorial predictions to the experimental data from a small discrete system has been shown in detail in Ref. Lepek_2021 .
Another difference between the two frameworks is the definition of the gelation point. In the traditional Smoluchowski approach, gelation is defined by the mass deficiency that appears after a critical time and is attributed to an infinite cluster (a gel). The gel can be assumed to be either passive or active with respect to the sol fraction. The passive gel grows due to a finite mass flux toward infinite particle sizes, while the active gel grows, in addition, by consuming the sol particles. In turn, in the Marcus–Lushnikov approach, where finite coagulating systems are considered, the gel manifests itself as a narrow hump in the distribution of the average particle numbers over their masses (cf. Fig. 1 in Ref. Fronczak_2019 or Fig. 8b below). This hump appears after the critical time at macroscopically large mass $s\propto N$ and behaves like the active gel Lushnikov_2006_review . Nevertheless, the appearance of a hump at large cluster sizes cannot be taken as a unique indicator of gelation, since, due to the finite system size, this feature eventually emerges for all kernels within the Marcus–Lushnikov framework, regardless of whether they exhibit gelation in the Smoluchowski model (cf. Fig. 8b and Fig. 8d). This contrasts with the Smoluchowski formulation, where gelation is a distinctive phenomenon that occurs only for kernels with aggregation rates that grow rapidly with cluster size (e.g., $K=sg$).